4c): Many freeway off-ramps, as well as corners at racetracks, are b the corners at the Indianapolis Motor Speedway are banked by 9.20°. The uation for the fastest speed at which a race radius of each corner is 256 m. Derive an eq car could take one of these corners if the coefficient of staticfriction is μs = 0.550? toward enter of circle
Answer #1

Sol 1 may

hence maximum velocity can be given as

v_{max}=\sqrt{\frac{Rg(\mu_s+\tan\theta)}{1-\mu_s\tan\theta}}=\sqrt{\frac{256\times9.8\times(0.550+\tan9.20)}{1-0.550\times\tan9.20}}\\ v_{max}=44.28$ m/s$